# Hypergraph adjusted plus-minus

### v2

@@ 5.1 Data Processing
In this section, we apply PAPM, HAPM, and LAPM to NBA data from 2012-2022. The raw play-by-play data can be purchased from https://www.bigdataball.com/.
Our data processing pipeline is as follows. For each team, we begin by computing the PM data for each game. We then aggregate PM over each season, i.e. we compute $X$, $W$, and $Y$ where the rows of $X$ are the unique lineups, $W$ is the total time played together for a given lineup, and $Y$ is that lineup’s PM over the course of the season. We restricted our analysis to only include players with at least 10,000 seconds of playing time, which is about 2 minutes per game. For comparison, the minimum time for all-pro award eligibility was changed in 2023 to 65 games and 20 min/game, which is 78,000 seconds. For simplicity, we do not adjust for the COVID shortened seasons (2019-2020 and 2020-2021).
@@
in which each row of $X$ has exactly five entries of +1 and five entries of -1. By setting $Y$ as the PM with respect to the home team, the negative entries of $X$ ensure that $Y_{j}<0$ corresponds to an away lineup with a positive PM. We refer to these metrics at the league level with defense as APM and PAPM, and at the team level without defense as APM_team and PAPM_team.
@@ 5.2 Rank Correlation
For each year from 2012-2022, we compute the Spearman rank correlation between each ranking method and the given advanced metric. The results are presented in Figure 3. For methods calculated based on team data (APM_team, PAPM_team, HAPM, and LAPM), the fitted values/coefficients across all teams are combined and ranked, so that all correlations are being calculated at the league level. Higher positive correlations imply greater agreement with benchmark rankings, and the box-plots display the variability in the correlations across the years.
@@
For all of the ridge regression models (APM, PAPM, and HAPM), we chose the regularization parameter $\lambda$ that “gives the most regularized model such that the cross-validated error is within one standard error of the minimum,” which is the default choice for the glmnet package (Hastie et al., 2021). We explored other criterion for choosing $\lambda$, but this valued maximized the correlation for APM and PAPM. On the other hand, HAPM was robust to different choices of $\lambda$.
For each of the advanced metrics, other than WS.48, we see that HAPM is better (i.e. in more agreement with the advanced metrics) than APM. In contrast, LAPM is most highly correlated with WS.48. This is important because WS.48 considers offensive and defensive contributions (rather than PER, which is primarily a measure of offensive production and scoring efficiency), incorporates playing time, and is highly connected to the corresponding team’s overall performance and win record.
From these comparisons, we see that HAPM and LAPM align with advanced metrics better than APM despite neither explicitly incorporating defensive considerations nor being computed at the league level. Moreover, HAPM and LAPM offer simultaneous rankings for all generalized lineups and are more robust to choice of penalization parameter.
@@ 5.3 Case Study
Joel Embiid 2295 365 14.1 (1 - 1) 166.7, 1 (1 - 1)
James Harden 793 151 8.9 (2 - 3) 73.9, 4 (4 - 5)
Tyrese Maxey 2648 192 8.0 (2 - 4) 94.2, 2 (2 - 3)
Tobias Harris 2542 170 7.8 (3 - 4) 88.6, 3 (3 - 3)
Danny Green 1349 77 3.0 (5 - 6) 56.8, 7 (6 - 10)
Seth Curry 1565 112 2.4 (5- 7) 68.3, 5 (4 - 6)
Matisse Thybulle 1685 88 1.9 (6 - 8) 60.1, 6 (6 - 8)
Georges Niang 1742 69 1.7 (7 - 8) 55.4, 8 (7 - 9)
@@
We show the top eight individual players according to HAPM in Table 5. We provide their HAPM and LAPM along with the 50% confidence intervals for the implied rank. The full distribution of HAPM/LAPM values is given in Figure A1 in the appendix. Unsurprisingly, we see the top four players are Embiid, Harden, Maxey, and Tobias Harris, a former first-round draft pick. Importantly, despite only having a PM of 151, Harden is identified as the second best player by HAPM. Although there are positive rank correlations between our metrics and PM (0.98 and 0.94 for HAPM and LAPM, respectively), by plotting them against PM in Figure 4, we can identify players that may be over or under-valued based on PM alone.
@@
We record the top ten pairs according to HAPM in Table 6 and visualize them in Figure 5. The rank correlations with PM are 0.58 and 0.59 for HAPM and LAPM, respectively. Here we see that not only is Embiid a great individual player, but he contributes to the top four and seven of the top ten best pairings according to HAPM (and the top three and seven of the top ten best pairings according to LAPM). This is consistent with Embiid’s reputation as a player with the ability to pass out of a double team. Furthermore, despite only playing together for 600 total minutes, HAPM and LAPM recognize the successful pairing of Embiid and Harden. As we hoped, our metrics identify Harden as a good teammate, belonging to three of the top six pairs for both HAPM and LAPM.
Interestingly, we see a high ranking for the pairing of Embiid and Matisse Thybulle (9th best pair according to HAPM and 5th best pair according to LAPM) despite having a PM of only 82. This can be explained by the excellent defensive pairing of Embiid and Thybulle, which both of our methods are able to detect. Not surprisingly, the worst ranking pair according to HAPM (shown in Figure 4) is Deandre Jordan and Paul Reed, who are both big men with overlapping skillsets. It is also interesting that the two best pairs according to PM (Embiid and Maxey & Embiid and Harris) have relatively low pair rankings (10 and 8 respectively) according to LAPM. This may suggest that the high PM of these pairings can be attributed to individual player success or the success of higher-order generalized lineups (e.g. Embiid and Harris play together in the 3rd, 9th, 13th and 15th best ranked trios according to LAPM).
@@
James Harden:Joel Embiid 193 22.9 (1 - 2) 67.1, 1 (2 - 9)
Joel Embiid:Tyrese Maxey 268 22.0 (1 - 3) 54.5, 10 (20 - 26)
Joel Embiid:Tobias Harris 256 21.9 (2 - 3) 55.6, 8 (17 - 23)
Danny Green:Joel Embiid 171 17.0 (5 - 8) 65.9, 2 (2 - 9)
James Harden:Tyrese Maxey 144 16.8 (4 - 10) 59.4, 6 (6 - 24)
James Harden:Tobias Harris 154 16.7 (4 - 11) 61.2, 4 (4 - 20)
Joel Embiid:Seth Curry 131 16.4 (6 - 9) 64.5, 3 (3 - 10)
Joel Embiid:Matisse Thybulle 82 15.9 (6 - 10) 59.8, 5 (6 - 18)
Tobias Harris:Tyrese Maxey 166 15.8 (5 - 13) 52.7, 17 (31 - 38)
Georges Niang:Joel Embiid 133 15.7 (7 - 10) 53.5, 12 (26 - 33)
@@
An immediate takeaway is the central role of Embiid, as he is part of 12 of the best 15 pairs and 14 of the 15 best trios according to HAPM. In both visualizations, he is at the center of the network. Exploratory analyses for other teams and other years reveals that this is very common; the superstars are the most central nodes in the visualizations (also shown for the Boston Celtics in Figure A3 in the appendix). Further investigation reveals that pair nodes that are larger than one or both of the player nodes (Embiid and Danny Green; Embiid and Thybulle; Embiid and Seth Curry; Embiid and Shake Milton) may represent possible synergies, whereas pair nodes that are smaller than the player nodes (Embidd and Isaiah Joe; Harris and Maxey) might indicate inefficient pairings.
@@ 6 Discussion
In this paper, we propose a new framework for ranking individual players using box-score plus-minus data, while simultaneously ranking all genearlized lineups. We show that our methods, HAPM and LAPM, are similarly or more correlated with advanced metrics for ranking individual players than ordinary APM. We also provide a case study on the Philadelphia 76ers in 2021-22 that seemingly identifies the best individuals, pairs, and trios. Lastly, we illustrate how our methods can easily be visualized in interpretable lineup networks.
@@
There are several directions for future research. Here, we presented two new metrics, LAPM and HAPM, for ranking generalized lineups. While we compared their results and stated the theoretical differences in the two approaches, we did not attempt to understand the cause of the metrics (dis)agreement. This may be a case-by-case exercise that could shed light on the inner workings of any one team. Furthermore, while it is straightforward to extend HAPM and LAPM to other APM variants that account for other non box-score statistics (Kubatko et al., 2007), it is less obvious how to incorporate defense for HAPM and LAPM for a league analysis rather than at the team level.

### v3

@@ 5.1 Data Processing
In this section, we apply PAPM, HAPM, and LAPM to NBA data from 2012-2022. The raw play-by-play data can be purchased from https://www.bigdataball.com/. Our code can be found at the GitHub repository njosephs/HAPM.
Our data processing pipeline is as follows. For each team, we begin by computing the raw plus-minus (PM) data for each game. We then aggregate PM over each season, i.e. we compute $X$, $W$, and $Y$ where the rows of $X$ are the unique lineups, $W$ is the average time per game played together for a given lineup, and $Y$ is that lineup’s PM over the course of the season. We restricted our analysis to only include players with at least 10,000 seconds of playing time, which is about 2 minutes per game. For comparison, the minimum time for all-pro award eligibility was changed in 2023 to 65 games and 20 min/game, which is 78,000 seconds. For simplicity, we do not adjust for the COVID shortened seasons (2019-2020 and 2020-2021), which, given the smaller sample size, will result in smaller between player variance for the APM statistics at the league level.
@@
in which each row of $X$ has exactly five entries of +1 and five entries of -1. By setting $Y$ as the PM with respect to the home team, the negative entries of $X$ ensure that $Y_{j}<0$ corresponds to an away lineup with a positive PM. We refer to these metrics at the league level with defense as APM_league and PAPM_league, and at the team level without defense as APM and PAPM, respectively.
Remark.
Our methods, HAPM and LAPM, are calculated at the team level and do not incorporate the opposing players, as this would require some notion of “negative edges.” Consequently, the responses for HAPM and LAPM (as well as APM and PAPM) represent the season-level net points for a generalized lineup, without considering the opposing players on the court. In contrast, the responses for APM_league and PAPM_league account for both the season-level net points of a generalized lineup and the specific set of opponents. Despite this distinction, we can still make league-level comparisons by aggregating fitted values across teams, as described in Section 5.2.
For all of the ridge regression models (APM, PAPM, and HAPM), we chose the regularization parameter $\lambda$ that minimizes the cross-validation error, which is one of the default choice for the glmnet package (Hastie et al., 2021).
@@ 5.2 Rank Correlation
For each year from 2012-2022, we compute the Spearman rank correlation between each ranking method and the given advanced metric. The results are presented in Figure 3. For methods calculated based on team data (APM, PAPM, HAPM, and LAPM), the fitted values/coefficients across all teams are combined and ranked, so that all correlations are being calculated at the league level. Higher positive correlations imply greater agreement with benchmark rankings, and the box-plots display the variability in the correlations across the years.
In Figure 3, we see that HAPM is better than APM, i.e. in more agreement with, for BPM, PER, and VORP. In contrast, LAPM is most highly correlated with WS.48. This is important because WS.48 considers offensive and defensive contributions (rather than PER, which is primarily a measure of offensive production and scoring efficiency), incorporates playing time, and is highly connected to the corresponding team’s overall performance and win record.
We can also see what is lost when moving away from APM_league, which is modeled at the league level and adjusts for defense. Moving left to right among the methods, we see that PAPM_league is generally better than APM_league for ranking individuals, which shows that adding interaction terms can improve the model’s performance. However, because including additional pair columns in the design matrix requires greater regularization (i.e. shrinkage of all of the coefficients), the improvement is not immense. We also see that computing APM at the team level is worse than at league level, which makes the performance of HAPM and LAPM even more noteworthy.
@@
From these comparisons, we see that HAPM and LAPM align with advanced metrics as well as APM_league despite neither explicitly incorporating defensive considerations nor being computed at the league level. Moreover, HAPM and LAPM offer simultaneous rankings for all generalized lineups and are more robust to choice of penalization parameter.
5.3 Predictive Accuracy
In addition to looking at the agreement between our methods and several advanced analytics, we can see how predictive our metrics are of future player/pair/generalized lineup performance. In Figure 4 (left) we display the year-over-year rank correlations of the APM methods for individual players. For years $t$ and $t+1$, we consider only the players that qualified for inclusion in both years, but we allow those players that switch teams. We see that HAPM and LAPM have higher correlations than the other APM variants, which suggests HAPM and LAPM are more stable metrics. However, we note that the rank correlation values are not particularly high, so there is clearly change happening in the rankings over time, which we should expect.
To evaluate within-season prediction capability, we split each season into two halves and assess how well the metrics predict data from the second half of the season when fit on data from the first half of the season. In Figure 4 (middle), we see that LAPM predictions have the highest rank correlation to the observed, second-half of season, plus-minus values. However, this evaluation interprets PM as the ground truth, which is somewhat problematic given our starting argument is that PM is not a complete reflection of an individual player’s talent. Therefore, we also include in the appendix (Figure A1) the rank correlation between halves of the season for each metric. This again suggests that HAPM and LAPM are the more stable metrics over time. That is, incorporating all generalized lineups in the modeling methodology (as HAPM and LAPM do) appears to help player ranking predictions.
We perform similar evaluations on the ranked pairs. In addition to PAPM, HAPM, and LAPM, we consider the two naive approaches mentioned in the introduction. In particular, we let $\Sigma$APM and PM denote the methods that sum the APM of all individuals and the cumulative PM, respectively, for the generalized lineup of interest (in this case, the pairs). Again, we see that LAPM is the most predictive method with HAPM also performing well. Interestingly, PM is very predictive for pairs, but as we noted, this evaluation method considers PM as the ground truth, so if there are dependencies in, for example, substitution patterns, then the second-half PM values will likely also be affected, which may be artificially inflating the correlation. As above, we include the correlation between season halves for each metric in Figure A1.
Figure 4: Spearman rank correlation for each method year-over-year (left), within year to observed PM for individuals (middle), and within year to observed PM for pairs (right) across the years 2012-2022.
@@ 5.4 Case Study
We show the top eight individual players according to HAPM in Table 5. We provide their HAPM and LAPM along with the 50% confidence intervals for the implied rank. The full distribution of HAPM/LAPM values is given in Figure A2 in the appendix. Unsurprisingly, we see the top four players are Embiid, Harden, Maxey, and Tobias Harris, a former first-round draft pick. Importantly, despite only having a PM of 151, Harden is identified as the second best player by HAPM. Although there are positive rank correlations between our metrics and PM (0.68 and 0.85 for HAPM and LAPM, respectively), by plotting them against PM in Figure 5, we can identify players that may be over or under-valued based on PM alone.
@@
Joel Embiid 2295 365 75.0 (1 - 1) 166.7, 1 (1 - 1)
James Harden 793 151 38.8 (2 - 3) 73.9, 4 (4 - 5)
Tyrese Maxey 2648 192 27.8 (3 - 4) 94.2, 2 (2 - 3)
Tobias Harris 2542 170 25.6 (3 - 4) 88.6, 3 (3 - 3)
George Niang 1742 69 0.8 (5 - 8) 55.4, 8 (7 - 9)
Danny Green 1349 77 -0.1 (5 - 9) 56.8, 7 (6 - 10)
Isaiah Joe 609 23 -2.2 (6 - 8) 48.4, (8 - 12)
Charles Bassey 168 -34 -4.3 (7 - 10) 45.1, (9 - 14)
@@
We record the top ten pairs according to HAPM in Table 6 and visualize them in Figure 6. The rank correlations with PM are 0.61 and 0.46 for HAPM and LAPM, respectively. Here we see that not only is Embiid a great individual player, but he contributes to the top nine pairings according to HAPM (and seven of the top ten best pairings according to LAPM). This is consistent with Embiid’s reputation as a player with the ability to pass out of a double team. Furthermore, despite only playing together for 600 total minutes, both HAPM and LAPM recognize the successful pairing of Embiid and Harden. As we hoped, our metrics identify Harden as a good teammate, belonging to three of the top six pairs for both HAPM and LAPM.
@@
James Harden:Joel Embiid 193 113.8 (1 - 2) 67.1, 1 (2 - 9)
Joel Embiid:Tyrese Maxey 268 102.8 (1 - 3) 54.5, 10 (20 - 26)
Joel Embiid:Tobias Harris 256 100.6 (2 - 4) 55.6, 8 (17 - 23)
Georges Niang:Joel Embiid 133 75.8 (5 - 9) 53.5, 12 (26 - 33)
Danny Green:Joel Embiid 171 74.9 (5 - 9) 65.9, 2 (2 - 9)
Isaiah Joe:Joel Embiid 25 72.8 (6 - 9) 53.1, 15 (13 - 78)
Andre Drummond:Joel Embiid 5 68.0 (7 - 14) 53.1, 16 (13 - 79)
Joel Embiid:Seth Curry 131 67.7 (8 - 12) 64.5, 3 (3 - 10)
Furkan Korkmaz:Joel Embiid 108 67.4 (8 - 13) 54.7 (19 - 28)
James Harden:Tyrese Maxey 144 66.5 (4 - 13) 59.4, 6 (6 - 24)
@@
Interestingly, we see high rankings for the pairings of Embiid and Andre Drummond (9th best pair according to HAPM and 16th best pair according to LAPM, out of 96 qualifying pairs) and Embiid and Matisse Thybulle (13th/5th according to HAPM/LAPM) despite having PMs of only 5 and 82, respectively. This can be explained by the excellent defensive pairing of Embiid with Drummond and Thybulle, which both of our methods are able to detect. Not surprisingly, one of the worst ranking pair according to HAPM (shown in Figure 6) is Deandre Jordan and Paul Reed, who are both big men with overlapping skillsets. It is also interesting that the two best pairs according to PM (Embiid and Maxey & Embiid and Harris) have relatively low pair rankings (10th and 8th, respectively) according to LAPM. This may suggest that the high PM of these pairings can be attributed to individual player success or the success of higher-order generalized lineups (e.g. Embiid and Harris play together in the 3rd, 9th, 13th and 14th best ranked trios according to LAPM).
@@
An immediate takeaway is the central role of Embiid, as he is part of 12 of the best 15 pairs and all of the 15 best trios according to HAPM. In both visualizations, he is at the center of the network. Exploratory analyses for other teams and other years reveals that this is very common; the superstars are the most central nodes in the visualizations (also shown for the Boston Celtics in Figure A4 in the appendix). Further investigation reveals that pair nodes that are larger than one or both of the player nodes (Embiid and Danny Green; Embiid and Thybulle; Embiid and Seth Curry; Embiid and Shake Milton) may represent possible synergies, whereas pair nodes that are smaller than the player nodes (Embidd and Isaiah Joe; Harris and Maxey) might indicate inefficient pairings.
@@ 6 Discussion
In this paper, we propose a new framework for ranking individual players using plus-minus data, while simultaneously ranking all genearlized lineups. We show that our methods, HAPM and LAPM, are similarly or more correlated with advanced metrics for ranking individual players than ordinary APM. We also provide a case study on the Philadelphia 76ers in 2021-22 that seemingly identifies the best individuals, pairs, and trios. In terms of utility, coaches can identify under- and over performing player combinations mid-season, adjust their lineups accordingly, and evaluate the impact on team performance. Moreover, network visualizations illustrate these player interactions, directly highlighting which combinations drive success or need improvement.
@@
There are several directions for future research. Here, we presented two new metrics, LAPM and HAPM, for ranking generalized lineups. While we compared their results and stated the theoretical differences in the two approaches, we did not attempt to understand the cause of the metrics (dis)agreement. This may be a case-by-case exercise that could shed light on the inner workings of any one team. Furthermore, while it is straightforward to extend HAPM and LAPM to other APM variants that account for other non box-score statistics (Kubatko et al., 2007), it is less obvious how to incorporate defense for HAPM and LAPM. Our work demonstrates that HAPM and LAPM can achieve comparable results to metrics calculated at the league level even without incorporating defense, but this is a natural future direction.
@@ Appendix B Appendix
Figure A1: Spearman rank correlation for each method within year to itself for individuals (left) and pairs (right) across the years 2012-2022.
Appendix C Appendix